OpenStudy (anonymous):

Order the following from least to greatese; 6.5, 6/5, 600%, 1/6, -60, 1/5 Please show work I dont know how to trun fractions and percents into decimals

OpenStudy (anonymous):

Got a calculator? if so fractions to decimals is easy 1/6 = 1 divided by 6 = 0.1666666666666667. percentage is easy too 1% = 1/100 or 0.01, 600% = 600/100 = 6